1. A survey of a sample of business students resulted in the following information regarding the genders of the individuals and their selected major.

Selected Major

Gender Management Marketing Others Total
Male 40 10 30 80
Female 30 20 70 120
Total 70 30 100 200

a. What is the probability of selecting an individual who is majoring in Marketing?
b. What is the probability of selecting an individual who is majoring in Management, given that the person is female?
c. Given that a person is male, what is the probability that he is majoring in Management?
d. What is the probability of selecting a male individual?
1
Expert's answer
2020-04-24T20:37:20-0400

a. Probability of selecting an individual who is majoring in Marketing: P=30/200=0.15.

b. Probability of selecting an individual who is majoring in Management, given that the person is female: P=30/120=0.25.

c. Given that a person is male, probability that he is majoring in Management: P=40/80=0.5.

d. Probability of selecting a male individual: P=80/200=0.4